Please help with the following:
4/5 x 1 1/6 =
4(-2)² + 8(-2) + 3(-2) + 6 =
-11(3/2)(.12(-12)) =
(1/3)9³ =
Thank you.

Answer by stanbon(75887)

4/5 x 1 1/6 = (4/5)*(7/6) = (2/3)(7/5) = 14/15
-------------------------------------------------------
4(-2)² + 8(-2) + 3(-2) + 6 = -8 + -16 + -6 = -30
---------------------------------------------------
-11(3/2)(.12(-12)) = (-33/2)(-1.44) = (-16.5)(-1.44) = 23.76
---------------------------------------------------
(1/3)9³ = (1/3)(3^6) = 3*5 = 243
=====================================
Cheers,
Stan H.
=============